OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (kymber):

The slope of parallel lines are the same, so plug in the point & slope (1) into y = mx + b & solve for b

OpenStudy (anonymous):

B and D option are same and correct. D. y = x + 6 for x =4 and y =6 only this equation holds good. Use common sense !